An airplane moves 100 m/s as it travels around a v
Answer #1

7). Velocity of plane, v= 100m/s

Radius, r= 1 km

Mass of pilot, M= 50 kg

Centripetal Force acting towards centre( vertically upwards direction when plane is at bottom) on pilot is,

F_c= Mv^2/r

  F_c= (50)(100)^2/(10^3)= 500N

Weight of pilot acting in downward, W=Mg= 50 \times 9.8= 490 N

Thus Net force acting on pilot, F_N=F_c-W=500- 490 = 10 N   ( towards the centre of loop)

8). Mass of block, m= 4 kg

Spring Constant, k= 200 N/m

Velocity of block at equilibrium position, v= 4 m/s

Let block initially compressed the spring to a distance x

thus Velocity gained by block will be due to the energy provided by block. There will be complete conversion of energy in this due to frictionless surface.

Hence K.E_S= K.E_b

  (1/2)kx^2= (1/2)mv^2

  x= \sqrt{mv^2/k}= \sqrt{ 4\times 4^2/200}= \sqrt{0.32}= 0.56 m

Now force exerted by spring on block is, Fs= kx

Also by newton's laws of motion, Fb= ma

Equating both forces as force on block is due to spring only.

hence ma=kx\Rightarrow a= kx/m

a= (200)(0.565)/4= 28.28 m/s^2

Now at equilibrium position, initial velocity, u= 4m/s

acceleration, a= 28.28 m/s2

Distance, S= 20cm= 0.2m

let final velocity at this much distance be v

thus by using equations of motion,

v^2-u^2=2aS

v^2-(4)^2=2(28.28)(0.2)

  v=\sqrt{11.31+16}= \sqrt{27.31}= 5.22 m/s  ( ANS)